# Theoretical analysis of performance limitation of computational refocusing in optical coherence tomography

### v1

@@ 5.1 Impact of lateral phase modulation by off-axis reference and BM-scan
5.2 Impact of multiple-scatter interactions
Note that the phase of the OCT signal can be modulated if multiple scatterers exist within a coherence volume (i.e., resolution volume). However, this modulation does not affect our Nyquist criterion for the following reasons.
In Section 5.1, we derived the criteria from the perspective of phase difference between adjacent pixels. However, in this analyses, we first split the PSF into the real Gaussian envelope and a phase-only function. Therefore, in reality, this phase analysis is not really an analysis of the phase difference but is an analysis of the local Nyquist frequency of complex PSF. In the complex signals, the contributions from multiple scatterers are linearly superposed, and this superposition does not change (i.e., broaden or narrow) the spatial frequency spectrum of the OCT image.
In summary, the presence of multiple scatterers in a coherence volume may lead to modulation of the OCT phase, but it does not alter the Nyquist criteria. This is not like the phase modulation case that was discussed in Section 5.1.
@@ 5.3 Limitations and solutions
The lack of confocality makes spatially-coherent FFOCT method advantageous for computational refocusing, but it can also cause image degradation because of the multiple-scattering (MS) signals. This problem can be resolved using methods that combine hardware modification with signal processing. For example, STOC-T, which is a variation of spatially-coherent FFOCT, overcomes the MS-signal related image degradation by using wavefront modulation and subsequent incoherence image averaging[30] based on temporal modulation. Multi-focus averaging (MFA) methods [36, 37, 38] represent a combination of sequential OCT image acquisition with focus position modulation using an electrically tunable lens with subsequent computational refocusing and complex averaging. Although these methods were demonstrated with respect to point-scanning OCT, they can also be applied to spatially-coherent FFOCT.
Another factor thats affect the imaging depth but is not considered in our analyses is signal attenuation caused by scattering and absorption characteristics of the sample. Specifically, even the measurement fulfills the Nyquist criterion and the confocality limit criterion, the signals cannot be observed if the signal attenuation is too high. High signal attenuation also affects the confocal function, although we assumed a naïve confocal function in our analyses (Eq. 24). Some studies are dedicated to define and/or measure light attenuation caused by sample sacattering and absorption in confocal functions [39, 40], and these modified confocal functions can be applied to make our analyses more accurate.
Another limitation in our analyses involves the usage of approximations in the pupil and spot descriptions. For point-scanning OCT, the modeling was based on the paraxial Gaussian model [20], which tacitly assumes that the lenses are aplanatic, and this is not an accurate approximation for very high NA cases. For spatially-coherent FFOCT, we approximated the collection spot using a Gaussian spot, whereas the spot is an Airy disc pattern in reality (see Section 2.2.2). In other words, we approximated the collection pupil using a Gaussian pupil, whereas it is a cylinder function with a clear cut-off frequency. These approximations are reasonable for most of the realistic cases, but some modification may be required to apply the analyses to very high NA cases.
In section 2.2.2, the illumination spot of the spatially-coherent FFOCT is a constant spot. It should be noted here that, although the illumination field (i.e., the illumination spot) does not extend infinitely largely, it can be reasonably considered as a constant as far as the illumination field is sufficiently large.
@@
5.4 Spatially incoherence full-field OCT
@@ 5.5 Conclusion
In this paper, a theoretical consideration of the limitations of holographic refocusing has been presented, and two types of criteria, i.e., the Nyquist criterion and confocality limit criterion, were derived. Specifically, point-scanning OCT and spatially-coherent FFOCT methods were modeled using a dual pupil-based formulation to derive their Nyquist criterion. The Nyquist criterion give the required sampling densities for holographic refocusing, and can be summarized as follows: “the defocus is correctable regardless of the defocus amount as long as the lateral pixel density (i.e., sampling density) is lower than 55.5% of the in-air lateral resolution.”
@@
Unlike spatially-coherent FFOCT, the point-scanning OCT is also restricted by the confocality limit criterion. In summary, if the SNR is 40 dB or 20 dB, the signal becomes unobservable at a defocus distance of 100 times or 10 times of the Rayleigh range, respectively, even after refocusing.
These results implies that, as far as the appropriate Nyquist criterion are fulfilled, computational refocusing can work well, even for very large defocus distances. Because the practical limit of the imaging depth of OCT is dominated by the signal attenuation caused by the sample scattering and absorption, the amount of defocus that is correctable via computational refocusing is not a practical limiting factor for the imaging depth, especially in the spatially-coherent FFOCT case. This also indicates that spatially-coherent FFOCT is a particularly suitable technique for optical coherence microscopy.
Appendix
In this appendix, we derived Eq. 15 from Eq. 14 as described in Sec. 3.1.1.
For point-scanning OCT, because the radius of the PSF is defined as the $1/e$-radius of the amplitude, $w(z_{d})/\sqrt{2}$, the maximum absolute phase increment, thus becomes
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(z_{d};x)\right|\right)_{x}=\left|\frac{\sqrt{2}nk_{0}w(z_{d})}{R(z_{d})}\Delta x\right|.$ (28)
where $\mathrm{max}(\quad)_{x}$ represents the maximum over $x$, and this maximum is obtained at $x=w(z_{d})/\sqrt{2}$ for point-scanning OCT.
@@
Then, by substituting the phase curvature $R(z_{d})$(Eq. 5), beam radius $w(z_{d})$(Eq. 4),the in-focus beam radius $w_{0}$, the refractive index $n$, and the wavenumber $k_{0}$[in Sec. 2.3.1] into Eq. 14, the maximum absolute phase increments $\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(z_{d};x)\right|\right)_{x}$ become
@@
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(z_{d};x)\right|\right)_{x}=\left|\frac{\frac{2\sqrt{2}n\pi}{\lambda_{0}}\sqrt{1+(z_{d}/z_{R})^{2}}}{z_{d}\left[1+(z_{R}/z_{d})^{2}\right]}\Delta x\right|$ (29)
Here, we introduce a normalized defocus distance by taking the fraction of defocus distance over Rayleigh length, i.e., $\zeta_{d}=z_{d}/z_{R}$. Next, in the denominator, we can multiply by Rayleigh length $z_{R}$ and then divide by Rayleigh length $z_{R}$ at the same time, but we express the former of the beam properties $z_{R}=\frac{n^{2}w_{0}^{2}\pi}{\lambda_{0}}$ and the latter in the form of Rayleigh length $z_{R}$ itself to simplify the variables contained in the equation, and thus the maximum absolute phase increments $\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(z_{d};x)\right|\right)_{x}$ then become
@@
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(z_{d};x)\right|\right)_{x}=\left|\frac{\frac{2\sqrt{2}n\pi}{\lambda_{0}}\sqrt{1+\left(z_{d}/z_{R}\right)^{2}}}{z_{d}/z_{R}\left[1+(z_{R}/z_{d})^{2}\right]\frac{n^{2}w_{0}^{2}\pi}{\lambda_{0}}}\Delta x\right|=\left|\frac{2\sqrt{2}}{nw_{0}}\frac{1}{\sqrt{1+1/(z_{d}/z_{R})^{2}}}\Delta x\right|$ (30)
As a result, the maximum absolute phase increments $\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(z_{d};x)\right|\right)_{x}$ can be written as new variables of $\zeta_{d}$, and can be expressed as
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(\zeta_{d};x)\right|\right)_{x}=\left|\frac{2\sqrt{2}}{nw_{0}}\frac{1}{1+1/\zeta_{d}^{2}}\Delta x\right|.$ (31)
Therefore, Eq. 15 has been derived.
@@
Core Research for Evolutional Science and Technology (JPMJCR2105). Japan Society for the Promotion of Science (P23365, 23KF0186, 21H01836, 22F22355, 22KF0058, 22K04962, 24KJ0510). Chinese Scholarship Council (202106845011). National Natural Science Foundation of China (62005123). Natural Science Foundation of Jiangsu Province (BK20190455). Nanjing University of Science and Technology Independent Research Fund (30919011226).
Acknowledgments
@@
Makita, Yasuno: Nikon (F), Santec (F), Sky Technology(F), Panasonic (F), Topcon (F), Kao Corp.(F). Fukutake: Nikon (E).

### v3

@@ 5.2 Limitations and solutions
The lack of confocality makes spatially-coherent FFOCT method advantageous for computational refocusing, but it can also cause image degradation because of the multiple-scattering (MS) signals. This problem can be resolved using methods that combine hardware modification with signal processing. For example, STOC-T, which is a variation of spatially-coherent FFOCT, overcomes the MS-signal related image degradation by using wavefront modulation and subsequent incoherent image averaging[49]. Multi-focus averaging (MFA) methods [50, 51, 52] represent a combination of sequential OCT image acquisition with focus position modulation using an electrically tunable lens with subsequent computational refocusing and complex averaging. Although these methods were demonstrated with respect to point-scanning OCT, they can also be applied to spatially-coherent FFOCT.
Another factor thats affect the imaging depth but is not considered in our analyses is signal attenuation caused by scattering and absorption characteristics of the sample. Specifically, even the measurement fulfills the Nyquist criterion and the confocality-limit criterion, the signals cannot be observed if the signal attenuation is too high.
Although the optical property and the structure of the sample can affects the confocal function, this effect is not recapitulated by the naïve confocal function used in our analyses [Eq. (22)]. Some studies are dedicated to define and/or measure more accurate depth-intensity profile [53, 54], and they can be used to enhance the accuracy of our analysis.
Another limitation in our analyses involves the usage of approximations in the pupil and spot descriptions. For point-scanning OCT, the modeling was based on the paraxial Gaussian model [35], which tacitly assumes that the lenses are aplanatic, and this is not an accurate approximation for very high NA cases. For spatially-coherent FFOCT, we approximated the collection spot using a Gaussian spot (see Section 2.2.2). In other words, we approximated the collection pupil using a Gaussian pupil, whereas it is a cylinder function with a clear cut-off frequency in reality. These approximations are reasonable for most of the realistic cases, but some modification may be required to apply the analyses to very high NA cases.
In section 2.2.2, the illumination spot of the spatially-coherent FFOCT is a constant, i.e., the uniform illumination. It should be noted here that, although the illumination field (i.e., the illumination spot) does not extend infinitely largely, it can be reasonably considered as a constant as far as the illumination field is sufficiently large.
@@
The present theory is limited to a two-dimensional lateral analysis. The modification of the theory by incorporating spatial three-dimensional pupil-based imaging theory [55] or four-dimensional (i.e., space and time) imaging theory [31, 32] may improve the accuracy and applicability of our analyses.
5.3 Spatially incoherent full-field OCT
@@ 5.4 Conclusion
[…]
@@
Unlike spatially-coherent FFOCT, the point-scanning OCT is also restricted by the confocality-limit criterion. In summary, if the SNR is 40 dB or 20 dB, the signal becomes unobservable at a defocus distance of 100 times or 10 times of the Rayleigh range, respectively, even after refocusing.
In conclusion, the correctable defocus amount of the point-scanning OCT is highly limited by the confocality in practice, not only by the Nyquist criteria. On the other hand, for spatially-coherent FFOCT, computational refocusing can work well even for very large defocus distances, as far as the lateral sampling density and the resolution of the system were designed to fulfill the Nyquist criterion. This indicates that spatially-coherent FFOCT is a particularly suitable technique for optical coherence microscopy.
Appendix A Derivation of Eq. (15)
Eq. (15) was derived from Eq. (14) as follows.
@@
By substituting the phase curvature $R(z_{d})$ [Eq. (5)], beam radius $w(z_{d})$ [Eq. (4)],the in-focus beam radius $w_{0}$, the refractive index $n$, and the wavenumber $k_{0}$, Eq. (14) becomes
@@
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(z_{d};x)\right|\right)_{x}=\left|\frac{\frac{2\sqrt{2}n\pi}{\lambda_{0}}w_{0}\sqrt{1+(z_{d}/z_{R})^{2}}}{z_{d}\left[1+(z_{R}/z_{d})^{2}\right]}\Delta x\right|.$ (25)
We then multiply the denominator by Rayleigh length $z_{R}$ and divide it by Rayleigh length $z_{R}$ at the same time. But the former $z_{R}$ is expressed as $z_{R}=\frac{nw_{0}^{2}\pi}{\lambda_{0}}$, while the latter is directly expressed as $z_{R}$. Namely, the denominator is multiplied by $\left(nw_{0}^{2}\pi/\lambda_{0}\right)/z_{R}$ as
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(z_{d};x)\right|\right)_{x}=\left|\frac{2\sqrt{2}}{w_{0}}\frac{z_{R}\sqrt{1+(z_{d}/z_{R})^{2}}}{z_{d}\left[1+(z_{R}/z_{d})^{2}\right]}\Delta x\right|.$ (26)
This equation can be rewritten as
@@
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(z_{d};x)\right|\right)_{x}=\left|\frac{2\sqrt{2}}{w_{0}}\frac{\operatorname{sgn}(z_{d})}{\sqrt{1+1/(z_{d}/z_{R})^{2}}}\Delta x\right|.$ (27)
Eq. (15) is finally derived by introducing a normalized defocus distance, $\zeta_{d}=z_{d}/z_{R}$, into this equation.
@@
Core Research for Evolutional Science and Technology (JPMJCR2105); Japan Society for the Promotion of Science (23KF0186, 21H01836, 22F22355, 22KF0058, 22K04962, 24KJ0510); Chinese Scholarship Council (202106845011); National Natural Science Foundation of China (62005123); Natural Science Foundation of Jiangsu Province (BK20190455).
@@
Makita, Yasuno: Nikon (F), Nidek (F), Santec (F), Sky Technology(F), Panasonic (F), Topcon (F), Kao Corp.(F). Fukutake: Nikon (E).